Consider the following scenario involving a wheat farmer: The wheat farmer sells two futures contracts (each contract =10000 bushels of wheat) at a price of $3.50 per bushel of wheat to help hedge wheat prices. At the time the contract expires the spot price of wheat is $3.00 per bushel. The wheat farmer harvested 25000 bushels of wheat and delivers the previously sold futures contracts. we wheat farmer sells two futures contracts. What is the total cashflow the farmer van expect from the sale of their wheat?

75000

90000

80000

70000

85000
Answer #1

ANSWER DOWN BELOW. FEEL FREE TO ASK ANY DOUBTS. THUMBS UP PLEASE.

Contract for 20,000 bushels hedged at 3.50

Total production = 25,000 bushels.

Unhedged = 25,000-20,000 = 5,000 bushels, which will sold at market price.

Total cash inflow = (20,000*3.50)+(5000*3) =85,000

Answer: 85,000